Company ABC purchases equipment for $500,000 and estimates its salvage value to be $10,000 with a...

  1. Company ABC purchases equipment for $500,000 and estimates its salvage value to be $10,000 with a useful life of 10 years. After 3 years, the company realizes that the equipment will only last a total of 6 years with a salvage value of $20,000. What is the depreciation charge for year 4 (assume SL method)? (round to nearest dollar)
Answer #1

Depreciation expense per year = (Cost - Salvage value)/Useful life

= (500,000 - 10,000)/10 years = 49,000 per year

Book value after 3 years

= 500,000 - (49,000*3 years) = 353,000

Now useful life is revised to 6 years. 3 years have already passed. So, remaining life = 6 years - 3 years = 3 years.

Depreciation for year 4

= (353,000 - 20,000) / 3 years

= 111,000